OpenStudy (vera_ewing):

FAN AND MEDAL 4. Graph the function: y = x^2 -9. Identify and explain the key features of your graph. Explain how to find them algebraically. 5. Using your graph from question 4, explain whether the average rate of change is increasing or decreasing, from left to right. Justify your observations by comparing the slopes calculated between at least three different pairs of points.

OpenStudy (vera_ewing):

@dan815 When you're done please help me!

OpenStudy (dan815):

do u know how y=x^2 graph looks

OpenStudy (vera_ewing):

Yes, I just need help explainng the key features of the graph and explaining how to find them algebraically

OpenStudy (dan815):

y=2x^2 is similiar to y=x^2 but its 2 times more compressed horizontally as y is increasing twice as much

OpenStudy (dan815):

u can see this is true by dividing both of them suppose y1=2x^2 y2=x^2 y1/y2=2

OpenStudy (dan815):